# Sampling by averaging: A multiscale approach to score estimation

### v1

@@ B.1.2 Underdamped system
Schedule function $\lambda_{t}$, value for $\lambda_{\delta}$ and $\tilde{\lambda}$, friction coefficient $\Gamma$, mass parameter $M$, number of sampling steps $L$, time discretisation $0=T_{0}<\dots<T_{L}=1/\kappa$, step size $h_{l}=T_{l+1}-T_{l}$. Constants for SROCK step from (14), (15).
@@
$\quad\quad\quad\quad V_{l}^{\prime}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}+\sqrt{h_{l}\Gamma}\xi_{l}^{(1)}$
$\quad\quad\quad\quad V_{l}^{\prime\prime}=\begin{cases}V_{l}^{\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l}-Y_{l}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ V_{l}^{\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
@@
$\quad\quad\quad\quad X_{l+1}=X_{l}+h_{l}M^{-1}V_{l}^{\prime\prime}$
@@
$\quad\quad\quad\quad\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,0}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,0}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,i-1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,i-1}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l}^{(2)}\end{cases}$
@@
$\quad\quad\quad\quad V_{l}^{\prime\prime\prime}=\begin{cases}V_{l}^{\prime\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l+1}-Y_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ V_{l}^{\prime\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l+1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
$\quad\quad\quad\quad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$
@@
   $\quad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$ ,   end for return $(X_{L},V_{L})$
@@ B.2.2 Underdamped system
$f_{t}\left(x,\mathbb{E}_{Y|x,v}[Y]\right)=m_{3,t}\mathbb{E}_{Y|x,v}[Y]+\tilde{\Sigma}_{12,t}\tilde{\Sigma}_{11,t}^{-1}\left(x-m_{1,t}\mathbb{E}_{Y|x,v}[Y]\right),$
and
Using the derivations above, we define a multiscale SDE system that enables sampling from the target distribution via the reverse SDE—without requiring prior estimation of the denoiser $\mathbb{E}_{Y\mid x,v}[Y]$. We consider a fast process governed by an overdamped Langevin SDE targeting the conditional distribution of $Y$ given the current state $(\overleftarrow{X}_{t},\overleftarrow{V}_{t})$, which takes the form
@@
Inspired by Dockhorn et al. [2022b], we propose a novel discretisation scheme which leverages symmetric splitting techniques [Leimkuhler and Matthews, 2013] justified by the symmetric Trotter splitting and the Baker–Campbell–Hausdorff formula [Trotter, 1959; Strang, 1968; Tuckerman, 2010]. The $l$-th iteration of the proposed numerical scheme consists of the following steps, with $h_{l}$ denoting the step size.
@@
Friction coefficient $\Gamma$, $M=\Gamma^{2}/4$, number of sampling steps $L$, time discretisation $0=T_{0}<\dots<T_{L}=T$, step size $h_{l}=T_{l+1}-T_{l}$. Constants for SROCK step from (14), (15). Functions $\hat{m}_{t}(\cdot,\cdot)$ and $\hat{\Sigma}_{t}$ from (23).
@@
$\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\nabla V_{\pi}\left({K_{l,0}}\right)+\nabla_{Y}\log\rho_{T-T_{l}}(X_{l+1/2},V_{l+1/2}^{\prime}|{K_{l,0}})\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\nabla V_{\pi}\left({K_{l,i-1}}\right)+\nabla_{Y}\log\rho_{T-T_{l}}(X_{l+1/2},V_{l+1/2}^{\prime}|{K_{l,i-1}})\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l},\quad\xi_{l}\sim\mathcal{N}(0,I)\end{cases}$
$(X_{l+1/2},V_{l+1/2})\sim\mathcal{N}(\hat{m}_{h_{l}/2}(X_{l+1},V_{l+1/2}^{\prime}),\hat{\Sigma}_{h_{l}/2}$
@@ C.1 Convergence of the slow-fast system to the averaged dynamics
We analyse the first two cases, as standard Langevin dynamics is known to converge to its invariant distribution when initialised with a warm start [Wu et al., 2022; Chewi et al., 2021; Lee et al., 2021a]. See Appendix B.1 for further discussion.
@@
$\leq\begin{cases}2\Bigg{\|}\begin{pmatrix}M^{-1}(v_{1}-v_{2})\\ \mathds{1}_{\{0\leq\lambda_{\kappa t}<{\lambda}_{\delta}\}}\frac{1}{\sqrt{1-\lambda_{\delta}}}\left(\nabla\log\nu\left(\frac{x_{1}-y}{\sqrt{1-\lambda_{\delta}}}\right)-\nabla\log\nu\left(\frac{x_{2}-y}{\sqrt{1-\lambda_{\delta}}}\right)\right)-\Gamma M^{-1}(v_{1}-v_{2})\end{pmatrix}\\ \quad\quad\quad+\begin{pmatrix}0\\ \mathds{1}_{\{\lambda_{\delta}\leq\lambda_{\kappa t}<\tilde{\lambda}\}}\frac{1}{\sqrt{1-\lambda_{\kappa t}}}\left(\nabla\log\nu\left(\frac{x_{1}-y}{\sqrt{1-\lambda_{\kappa t}}}\right)-\nabla\log\nu\left(\frac{x_{2}-y}{\sqrt{1-\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\Bigg{\|}\|(x_{1},v_{1})-(x_{2},v_{2})\|&\text{Regime 1}\\ 0&\text{Regime 2}\end{cases}$
@@
$\leq\begin{cases}\Bigg{\|}\begin{pmatrix}0\\ \frac{\mathds{1}_{\{0\leq\lambda_{\kappa t}<{\lambda}_{\delta}\}}}{\sqrt{1-\lambda_{\delta}}}\left(\nabla\log\nu\left(\frac{x-y_{1}}{\sqrt{1-\lambda_{\delta}}}\right)-\nabla\log\nu\left(\frac{x-y_{2}}{\sqrt{1-\lambda_{\delta}}}\right)\right)\end{pmatrix}\\ \quad\quad\quad+\begin{pmatrix}0\\ \frac{\mathds{1}_{\{\lambda_{\delta}\leq\lambda_{\kappa t}<\tilde{\lambda}\}}}{\sqrt{1-\lambda_{\kappa t}}}\left(\nabla\log\nu\left(\frac{x-y_{1}}{\sqrt{1-\lambda_{\kappa t}}}\right)-\nabla\log\nu\left(\frac{x-y_{2}}{\sqrt{1-\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\Bigg{\|}&\text{Regime 1}\\ \left\|\begin{pmatrix}0\\ \frac{1}{\sqrt{\lambda_{\kappa t}}}\left(\nabla V_{\pi}\left(\frac{y_{1}}{\sqrt{\lambda_{\kappa t}}}\right)-\nabla V_{\pi}\left(\frac{y_{2}}{\sqrt{\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\right\|&\text{Regime 2}\end{cases}$
@@
$\quad+\left|\frac{m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}}{\sigma_{T-t}^{2}}-\frac{m_{3,T-s}-\tilde{\Sigma}_{12,T-s}\tilde{\Sigma}_{11,T-s}^{-1}m_{1,T-s}}{\sigma_{T-s}^{2}}\right|\|y\|\Bigg{]}$
@@
$\quad+\frac{4\Gamma}{\sigma_{T-t}^{2}}\left(m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}\right)\langle y,v\rangle$
@@
$\quad+\frac{2\Gamma}{\sigma_{T-t}^{2}}\left(m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}\right)\|y\|^{2}\leq C_{t}\|(x,v)\|^{2}+\lambda_{1}\|y\|^{2},$

### v2

@@ B.1.2 Underdamped system
Schedule function $\lambda_{t}$, value for $\lambda_{\delta}$ and $\tilde{\lambda}$, friction coefficient $\Gamma$ (scalar), mass parameter $M$ (scalar), number of sampling steps $L$, time discretisation $0=T_{0}<\dots<T_{L}=1/\kappa$, step size $h_{l}=T_{l+1}-T_{l}$. Constants for SROCK step from (14), (15).
@@
$\qquad\qquad\qquad\qquad V_{l}^{\prime}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}+\sqrt{h_{l}\Gamma}\xi_{l}^{(1)}$
$\qquad\qquad\qquad\qquad V_{l}^{\prime\prime}=\begin{cases}V_{l}^{\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l}-Y_{l}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ V_{l}^{\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
@@
$\qquad\qquad\qquad\qquad X_{l+1}=X_{l}+h_{l}M^{-1}V_{l}^{\prime\prime}$
@@
$\qquad\qquad\qquad\qquad\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,0}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,0}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\frac{1}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{K_{l,i-1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)+\frac{1}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{K_{l,i-1}-X_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l}^{(2)}\end{cases}$
@@
$\qquad\qquad\qquad\qquad V_{l}^{\prime\prime\prime}=\begin{cases}V_{l}^{\prime\prime}+\frac{h_{l}}{2\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\nabla\log\nu\left(\frac{X_{l+1}-Y_{l+1}}{\sqrt{1-\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}<\tilde{\lambda}$}\\ V_{l}^{\prime\prime}-\frac{h_{l}}{2\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\nabla V_{\pi}\left(\frac{Y_{l+1}}{\sqrt{\lambda_{\kappa T_{l}}^{\prime}}}\right)&\text{if $\lambda_{\kappa T_{l}}^{\prime}\geq\tilde{\lambda}$}\end{cases}$
$\qquad\qquad\qquad\qquad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$
@@
   $\quad V_{l+1}=\left(1-\frac{h_{l}}{2}\Gamma M^{-1}\right)V_{l}^{\prime\prime\prime}+\sqrt{h_{l}\Gamma}\xi_{l}^{(3)}$   end for return $(X_{L},V_{L})$
@@ B.2.2 Underdamped system
We note that the conditional distribution $\rho_{t}(v|x,y)$ remains Gaussian, with mean
$\tilde{m}_{t}(x,y)=m_{3,t}y+\tilde{\Sigma}_{12,t}\tilde{\Sigma}_{11,t}^{-1}\left(x-m_{1,t}y\right),$
which is linear in $y$ and covariance $\sigma_{t}^{2}I$, where
@@
$f_{t}\left(x,\mathbb{E}_{Y|x,v}[Y]\right)=m_{3,t}\mathbb{E}_{Y|x,v}[Y]+\tilde{\Sigma}_{12,t}\tilde{\Sigma}_{11,t}^{-1}\left(x-m_{1,t}\mathbb{E}_{Y|x,v}[Y]\right).$
Using the derivations above, we define a multiscale SDE system that enables sampling from the target distribution via the reverse SDE, without requiring prior estimation of the denoiser $\mathbb{E}_{Y\mid x,v}[Y]$. We consider a fast process governed by an overdamped Langevin SDE targeting the conditional distribution of $Y$ given the current state $(\overleftarrow{X}_{t},\overleftarrow{V}_{t})$, which takes the form
@@
Inspired by Dockhorn et al. [2022b], we propose a novel discretisation scheme which leverages symmetric splitting techniques [Leimkuhler and Matthews, 2013] justified by the symmetric Trotter splitting and the Baker–Campbell–Hausdorff formula [Trotter, 1959, Strang, 1968, Tuckerman, 2010]. The $l$-th iteration of the proposed numerical scheme consists of the following steps, with $h_{l}$ denoting the step size.
@@
Friction coefficient $\Gamma$ (scalar), $M=\Gamma^{2}/4$, number of sampling steps $L$, time discretisation $0=T_{0}<\dots<T_{L}=T$, step size $h_{l}=T_{l+1}-T_{l}$. Constants for SROCK step from (14), (15). Functions $\hat{m}_{t}(\cdot,\cdot)$ and $\hat{\Sigma}_{t}$ from (23).
@@
$\begin{cases}K_{l,0}&=Y_{l}\\ K_{l,1}&=K_{l,0}+\mu_{1}\frac{h_{l}}{\varepsilon}\left(-\nabla V_{\pi}\left({K_{l,0}}\right)+\nabla_{Y}\log\rho_{T-T_{l}}(X_{l+1/2},V_{l+1/2}^{\prime}|{K_{l,0}})\right)\\ &\vdots\\ K_{l,i}&=\mu_{i}\frac{h_{l}}{\varepsilon}\left(-\nabla V_{\pi}\left({K_{l,i-1}}\right)+\nabla_{Y}\log\rho_{T-T_{l}}(X_{l+1/2},V_{l+1/2}^{\prime}|{K_{l,i-1}})\right)+\nu_{i}K_{l,i-1}+\kappa_{i}K_{l,i-2}\\ &\vdots\\ Y_{l+1}&=K_{s}+\sqrt{\frac{2h}{\varepsilon}}\xi_{l},\qquad\xi_{l}\sim\mathcal{N}(0,I)\end{cases}$
$(X_{l+1/2},V_{l+1/2})\sim\mathcal{N}(\hat{m}_{h_{l}/2}(X_{l+1},V_{l+1/2}^{\prime}),\hat{\Sigma}_{h_{l}/2})$
@@ C.1 Convergence of the slow-fast system to the averaged dynamics
We analyse the first two cases, as standard Langevin dynamics is known to converge to its invariant distribution when initialised with a warm start [Wu et al., 2022, Chewi et al., 2021, Lee et al., 2021a]. See Appendix B.1 for further discussion.
@@
$\leq\begin{cases}2\Bigg\|\begin{pmatrix}M^{-1}(v_{1}-v_{2})\\ \mathds{1}_{\{0\leq\lambda_{\kappa t}<{\lambda}_{\delta}\}}\frac{1}{\sqrt{1-\lambda_{\delta}}}\left(\nabla\log\nu\left(\frac{x_{1}-y}{\sqrt{1-\lambda_{\delta}}}\right)-\nabla\log\nu\left(\frac{x_{2}-y}{\sqrt{1-\lambda_{\delta}}}\right)\right)-\Gamma M^{-1}(v_{1}-v_{2})\end{pmatrix}\\ \qquad\qquad\qquad+\begin{pmatrix}0\\ \mathds{1}_{\{\lambda_{\delta}\leq\lambda_{\kappa t}<\tilde{\lambda}\}}\frac{1}{\sqrt{1-\lambda_{\kappa t}}}\left(\nabla\log\nu\left(\frac{x_{1}-y}{\sqrt{1-\lambda_{\kappa t}}}\right)-\nabla\log\nu\left(\frac{x_{2}-y}{\sqrt{1-\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\Bigg\|\|(x_{1},v_{1})-(x_{2},v_{2})\|&\text{Regime 1}\\ 0&\text{Regime 2}\end{cases}$
@@
$\leq\begin{cases}\Bigg\|\begin{pmatrix}0\\ \frac{\mathds{1}_{\{0\leq\lambda_{\kappa t}<{\lambda}_{\delta}\}}}{\sqrt{1-\lambda_{\delta}}}\left(\nabla\log\nu\left(\frac{x-y_{1}}{\sqrt{1-\lambda_{\delta}}}\right)-\nabla\log\nu\left(\frac{x-y_{2}}{\sqrt{1-\lambda_{\delta}}}\right)\right)\end{pmatrix}\\ \qquad\qquad\qquad+\begin{pmatrix}0\\ \frac{\mathds{1}_{\{\lambda_{\delta}\leq\lambda_{\kappa t}<\tilde{\lambda}\}}}{\sqrt{1-\lambda_{\kappa t}}}\left(\nabla\log\nu\left(\frac{x-y_{1}}{\sqrt{1-\lambda_{\kappa t}}}\right)-\nabla\log\nu\left(\frac{x-y_{2}}{\sqrt{1-\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\Bigg\|&\text{Regime 1}\\ \left\|\begin{pmatrix}0\\ \frac{1}{\sqrt{\lambda_{\kappa t}}}\left(\nabla V_{\pi}\left(\frac{y_{1}}{\sqrt{\lambda_{\kappa t}}}\right)-\nabla V_{\pi}\left(\frac{y_{2}}{\sqrt{\lambda_{\kappa t}}}\right)\right)\end{pmatrix}\right\|&\text{Regime 2}\end{cases}$
@@
$\qquad+\left|\frac{m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}}{\sigma_{T-t}^{2}}-\frac{m_{3,T-s}-\tilde{\Sigma}_{12,T-s}\tilde{\Sigma}_{11,T-s}^{-1}m_{1,T-s}}{\sigma_{T-s}^{2}}\right|\|y\|\Bigg]$
@@
$\qquad+\frac{4\Gamma}{\sigma_{T-t}^{2}}\left(m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}\right)\langle y,v\rangle$
@@
$\qquad+\frac{2\Gamma}{\sigma_{T-t}^{2}}\left(m_{3,T-t}-\tilde{\Sigma}_{12,T-t}\tilde{\Sigma}_{11,T-t}^{-1}m_{1,T-t}\right)\|y\|^{2}\leq C_{t}\|(x,v)\|^{2}+\lambda_{1}\|y\|^{2},$